### 25 Mar 2026 v1

# Can LLMs Beat Classical Hyperparameter Optimization Algorithms? A Study on autoresearch

Abstract
The autoresearch repository enables an LLM agent to search for optimal hyperparameter configurations on an unconstrained search space by editing the training code directly. Given a fixed compute budget and constraints, we use autoresearch as a testbed to compare classical hyperparameter optimization (HPO) algorithms against LLM-based methods on tuning the hyperparameters of a small language model. Within a fixed hyperparameter search space, classical HPO methods such as CMA-ES and TPE consistently outperform LLM-based agents. However, an LLM agent that directly edits training source code in an unconstrained search space narrows the gap to classical methods substantially despite using only a self-hosted open-weight 27B model. Methods that avoid out-of-memory failures outperform those with higher search diversity, suggesting that reliability matters more than exploration breadth. While small and mid-sized LLMs struggle to track optimization state across trials, classical methods lack domain knowledge. To bridge this gap, we introduce Centaur, a hybrid that shares CMA-ES’s internal state, including mean vector, step-size, and covariance matrix, with an LLM [^1] Named after the mythological half-human, half-horse hybrid: our method merges LLM reasoning with classical optimization.. Centaur achieves the best result in our experiments, with its 0.8B variant outperforming the 27B variant, suggesting that a cheap LLM suffices when paired with a strong classical optimizer. The 0.8B model is insufficient for unconstrained code editing but sufficient for hybrid optimization, while scaling to 27B provides no advantage for fixed search space methods with the open-weight models tested. Code is available at https://github.com/ferreirafabio/autoresearch-automl.
1 Introduction
autoresearch (Karpathy, 2025a) demonstrated that an LLM agent can iteratively edit training code to improve a small language model ($\sim$}50M parameters), reaching a val_bpb of $\sim$}0.978 in less than 100 trials on a single H100. Building on that, Ziv (2025) showed that TPE, a classical AutoML method, with expert-designed hyperparameters can outperform Karpathy’s agent within a similar budget. Hyperparameter optimization (HPO) is a central component of AutoML (Feurer and Hutter, 2019; Bischl et al., 2023). These results raise two questions: (i) how do other classical HPO methods perform on this task? and (ii) can LLM-based HPO methods outperform classical ones?
To answer these questions, we benchmark 9 HPO methods, spanning 4 classical, 4 LLM-based, and 1 hybrid, on Karpathy’s autoresearch task. All methods operate under the same 24-hour GPU training budget with 3 seeds. To reduce human priors, we automatically extract 14 hyperparameters from the training script; while the ranges require some domain knowledge, the HP selection itself is automated, removing manual search space curation. Figure˜1 compares all 9 methods using 27B LLM variants against cumulative training wall-time, showing that classical methods find better configurations than LLM-based agents within the fixed search space. The exception is Karpathy Agent (Code), which directly edits training source code and is competitive with classical methods, even though the latter find a similar performing hyperparameter configuration ${\sim}4{\times}$ faster. Since all LLM methods use a self-hosted open-weight model (Qwen3.5-27B), stronger frontier models may narrow this gap further.
Beyond this comparison, we propose Centaur, a hybrid method that combines CMA-ES with an LLM by sharing the optimizer’s full internal state, including the mean vector $\bm{\mu}$, step-size $\sigma$, and covariance matrix $\mathbf{C}$. The hypothesis is that CMA-ES and LLMs have complementary strengths: CMA-ES learns the optimization landscape but lacks domain knowledge, while the LLM brings transformer training intuitions but, at the small and mid-sized scales we test (0.8B and 27B), struggles to track optimization state across trials reliably, e.g., LLM methods show OOM rates comparable to random search despite observing full trial history. This motivates pairing the LLM with a classical optimizer whose state can be shared explicitly. We chose CMA-ES because its internal state is particularly interpretable for LLM communication (see Section˜4).
In summary, we make the following contributions:
  • We benchmark 9 HPO methods on autoresearch (Karpathy, 2025a), supporting both fixed-HP and agentic code-editing optimization, under identical 24-hour budgets with 3 seeds.
  • We show that classical HPO outperforms LLM agents within a fixed search space, while an LLM agent editing code directly narrows the gap substantially despite using only a self-hosted open-weight model.
  • We introduce Centaur, a hybrid that shares CMA-ES’s full internal state with the LLM and achieves the best result in our experiments.
  • We analyze search diversity, OOM rates, and model scaling across all methods, and release per-trial LLM conversation logs.
Figure 1: Best Validation Bits-Per-Byte (mean $\pm$ std across 3 seeds) of HPO algorithms against cumulative training time. All methods receive the same 24-hour GPU training budget; LLM inference overhead is excluded. All LLM-based methods use Qwen3.5-27B as the LLM optimizer. Classical methods such as CMA-ES and TPE converge faster and to better final values than LLM-based methods. Centaur, our CMA-ES and LLM hybrid, achieves the best result in our experiments.
2 Related Work
Classical HPO spans a wide range of approaches (Bischl et al., 2023; Feurer and Hutter, 2019), from random search (Bergstra and Bengio, 2012) and Bayesian optimization with Gaussian process surrogates (Snoek et al., 2012) to sequential model-based approaches with random forests such as SMAC (Hutter et al., 2011), tree-structured Parzen estimators such as TPE (Bergstra et al., 2011), and evolution strategies such as CMA-ES (Hansen, 2016). We focus on single-task HPO methods without multi-fidelity or transfer to isolate each optimizer’s ability to learn the landscape from scratch. Methods such as Hyperband (Li et al., 2017), BOHB (Falkner et al., 2018), transfer HPO (Wistuba and Grabocka, 2021; Arango et al., 2024), and zero-shot approaches (Öztürk et al., 2022) could in principle be applied to this benchmark but are not in the scope of this study.
In line with the increasing interest in open-ended agentic discovery (Zhang et al., 2026; Novikov et al., 2025; Lange et al., 2026; Liu et al., 2024; Wang et al., 2026), recent work explores LLMs as components in HPO pipelines. While in principle, methods such as AlphaEvolve (Novikov et al., 2025) or ShinkaEvolve (Lange et al., 2026) can be tasked to optimize the objective in autoresearch, we focus only on methods tailored specifically for HPO. LLAMBO (Ye et al., 2024) uses an LLM as a surrogate model inside Bayesian optimization, replacing the Gaussian process with LLM-based performance predictions. SLLMBO (Mahammadli and Ertekin, 2024) integrates an LLM with TPE into a joint sampler. Zhang et al. (2023) prompt LLMs directly for HP suggestions. LLaMA-ES (Kramer, 2024) uses LLMs to tune CMA-ES’s own hyperparameters. Schwanke et al. (2026a) partition the search space into subregions via a bandit mechanism and use an LLM to propose candidates within each region. The authors further extend that to multi-objective optimization (Schwanke et al., 2026b). However, none of these methods share the classical optimizer’s full internal state with the LLM, which is the key idea behind Centaur.
Our work differs from prior work in four ways. First, we use autoresearch (Karpathy, 2025a) as a real-world benchmark that naturally accommodates both classical HPO within a fixed search space and agentic LLM optimization through direct code editing, enabling a head-to-head comparison under identical conditions. Second, we benchmark classical and LLM methods on this task following best practices for algorithm configuration (Eggensperger et al., 2019). Third, we automatically extract HPs from the training script via Abstract Syntax Tree (AST) parsing to control for human priors in the search space. Fourth, we introduce Centaur, a hybrid that explicitly passes CMA-ES’s mean, $\sigma$, and $\mathbf{C}$ to the LLM, enabling optimizer-informed suggestions rather than history-only LLM reasoning. In contrast, LLAMBO uses the LLM as a surrogate where the acquisition function decides, SLLMBO combines LLM and TPE proposals without exposing optimizer internals, and HOLLM (Schwanke et al., 2026a) uses spatial partitioning to constrain the LLM rather than inform it with optimizer state.
3 Experimental Setup
We describe the benchmark task and hardware, then the evaluation protocol and failure handling, followed by the LLM infrastructure, and finally the automated search space extraction.
We evaluate all methods on nanochat (Karpathy, 2025b), a small decoder-only transformer (Radford et al., 2019) trained on FineWeb (Penedo et al., 2024), optimizing validation bits-per-byte (val_bpb). Each trial ran for five minutes on a single NVIDIA H200 GPU (141 GB HBM3e).
We ran each method for 24 hours with three seeds and report results both against 24-hour wall-clock time and on trial scale (up to 300 trials). Methods with high OOM rates accumulated more trials due to fast failures (<30{<}30s per OOM) but show negligible improvement beyond trial 300. We report failed trials (OOM) as val_bpb=100.0{}=100.0, a finite penalty orders of magnitude worse than any valid result while remaining compatible with all surrogate models, so that optimizers learn to avoid infeasible regions.
All LLM-based methods use Qwen3.5 (Team, 2026) as the LLM optimizer, self-hosted via vLLM (Kwon et al., 2023) on the same GPU that trains the optimizee (the $\sim$}50M-parameter language model). To ensure equal resource allocation, we capped training VRAM to 80 GB for all methods, comparable to the H100 used by Karpathy (Karpathy, 2025a), and reserved the remaining memory for the vLLM server. We disabled Qwen3.5’s thinking mode and sampled with temperature 0.7, limiting outputs to 2048 tokens for fixed-HP methods and 16384 tokens for code editing. All prompt templates are available in our repository.
To extract hyperparameters from train.py, we used Abstract Syntax Tree (AST) parsing: we walk through parse tree and collect all top-level assignments whose variable name is in ALL_CAPS format and have a right-hand as a literal value (integer, float, or string). This yielded 14 HPs (13 continuous/integer and one categorical), listed in Table˜1 with their types, ranges, and defaults from Karpathy’s starting configuration. While the ranges require some domain knowledge, the hyperparameter selection itself is fully automated in this way, removing manual search space curation. Table˜2 summarizes the nine methods we evaluated across four classical, four LLM-based, and one hybrid approach.
Table 1: Search space: 14 HPs auto-extracted via AST parsing. Ranges are set manually. Defaults are Karpathy’s starting config. WINDOW_PATTERN controls the per-layer attention window: S = short (local) attention, L = long (full) attention.
HP Type Range Log Default
DEPTH int 4–24 8
ASPECT_RATIO int 32–128 64
HEAD_DIM int 64–256 128
DEVICE_BATCH_SIZE int 32–256 128
TOTAL_BATCH_SIZE int 65 536–2 097 152 524 288
EMBEDDING_LR float 0.01–2.0 0.6
UNEMBEDDING_LR float 0.0005–0.05 0.004
MATRIX_LR float 0.005–0.2 0.04
SCALAR_LR float 0.05–2.0 0.5
WEIGHT_DECAY float 0.0–0.5 0.2
WARMUP_RATIO float 0.0–0.3 0.0
WARMDOWN_RATIO float 0.1–0.8 0.5
FINAL_LR_FRAC float 0.0–0.2 0.0
WINDOW_PATTERN cat. {SSSL, SSLL, SLSL, LLLL, SSSS, LSSL} SSSL
Table 2: Methods evaluated. “Fixed” indicates the shared 14-HP search space; “Unconstrained” indicates direct source code editing.
Method Search Space Description
TPE Fixed Tree-structured Parzen Estimator via Optuna (Bergstra et al., 2011; Akiba et al., 2019)
CMA-ES Fixed Covariance Matrix Adaptation via Optuna’s CMA sampler (Hansen, 2016)
SMAC Fixed Random forest surrogate (Hutter et al., 2011)
Random Fixed Uniform random sampling
LLAMBO (Optuna) Fixed OptunaHub port: binary surrogate labels, random categorical sampling
LLAMBO (Paper) Fixed Reimplementation faithful to Ye et al. (2024): continuous labels, all HPs visible
Karpathy Agent (14 HPs) Fixed LLM sees trial history, suggests configs within fixed space
Karpathy Agent (Code) Unconstrained LLM directly edits train.py source code (Karpathy, 2025a)
Centaur Fixed CMA-ES shares internal state with LLM (Section˜4)
4 Centaur: CMA-ES Guided LLM Optimization
Centaur shares CMA-ES’s full internal state with the LLM on a fraction of trials. On every trial, CMA-ES proposes a candidate configuration from its multivariate Gaussian, parameterized by mean vector $\bm{\mu}$, step-size $\sigma$, and covariance matrix $\mathbf{C}$. On 30% of trials, the LLM receives CMA-ES’s proposal along with $\bm{\mu}$, $\sigma$, $\mathbf{C}$, the top-5 configurations, and the last 20 trials. The LLM may override the proposal, and in practice does so in nearly all cases: 100% of the time with the 27B model and 95% with the 0.8B model. Crucially, CMA-ES updates its internal state from all trial results, including those where the LLM overrode its proposal, so the optimizer continuously learns from the full trajectory.
We chose CMA-ES because its internal state is particularly interpretable for LLM communication: the mean is a concrete configuration, $\sigma$ is a single scalar, and $\mathbf{C}$ is a labeled matrix. In contrast, TPE maintains density estimators that are difficult to summarize in natural language, and GP-BO maintains a high-dimensional posterior. We provide the full procedure in Algorithm˜1.
Input: Search space $\mathcal{S}$, budget $T$, LLM ratio r=0.3r{=}0.3
1 Initialize CMA-ES, $\mathcal{H}\leftarrow\emptyset$;
2 for $t=1,\dots,T$ do
3 if LLM turn (with probability $r$) then
4      Extract $\bm{\mu},\sigma,\mathbf{C}$ from CMA-ES;
5    $\mathbf{x}\leftarrow$ LLM($\bm{\mu},\sigma,\mathbf{C}$, $\mathcal{H}$, $\mathcal{S}$);
6
7 else
8    $\mathbf{x}\leftarrow$ CMA-ES.Propose();
9
10   end if
11 $y\leftarrow$ Evaluate($\mathbf{x}$);
12   CMA-ES.Update($\mathbf{x},y$);
13 $\mathcal{H}\leftarrow\mathcal{H}\cup\{(\mathbf{x},y)\}$;
14
15 end for
Algorithm 1 Centaur
5 Results
[…]
5.1 Classical methods outperform LLMs in fixed search spaces
Figure˜1 shows convergence curves against cumulative training wall-time for all 27B methods. Within the fixed search space, classical HPO methods consistently outperformed pure LLM-based approaches. The top methods by mean best val_bpb are Centaur (0.9763), TPE (0.9768), SMAC (0.9778), CMA-ES (0.9785), and Karpathy Agent (Code) (0.9814). The gap to the best fixed-space LLM method, LLAMBO (Paper) at 0.9862, is substantial, and several pure LLM methods performed worse than random search, indicating that restricting LLMs to a fixed HP search space does not leverage their strengths.
OOM avoidance matters more than search diversity. Table˜3 shows that the top methods all have OOM rates at or below 16%, while the bottom four exceed 36%. Karpathy Agent (14 HPs) has the lowest diversity by all metrics (Table˜3), converging to a narrow region early. LLAMBO (Paper) is one of the most diverse methods yet underperforms classical methods due to its 48% OOM rate. These OOM rates also illustrate the state-tracking limitation of small and mid-sized LLMs: LLAMBO (Paper) and LLAMBO (Optuna) observe full trial history yet produce OOM rates (48% and 61%) comparable to random search (56%), suggesting they fail to learn which regions of the search space cause memory failures. In contrast, CMA-ES and TPE maintain explicit optimization state and keep OOM rates at 16% and 11%. CMA-ES explores broadly but efficiently: it covers 802 grid cells with only 16% OOM, showing that covariance adaptation learns which directions are safe. However, CMA-ES also has the highest variance among top methods (std 0.0036 vs 0.0019 for TPE), with its best seed achieving the single best result in the benchmark (0.9741) while its worst seed is mediocre (0.9829).
We report both LLAMBO variants because the OptunaHub implementation differs from the original paper in how it handles surrogate labels, categorical HPs, and failed trials, which substantially affects OOM rates and performance (see Section˜A.8 for details).
[…]
Method Trials Best val_bpb OOM% Spread Pairwise Step Cells
[…]
Figure 2: 0.8B vs 27B LLM optimizer comparison (wall-time). Solid: 27B, dashed: 0.8B. TPE and Random shown as classical references. The 0.8B model appears insufficient for unconstrained code editing but sufficient for hybrid optimization.
5.2 Unconstrained code editing is viable but requires model scale
Karpathy Agent (Code), which directly edits training source code rather than operating in the fixed search space, is the only pure LLM method competitive with classical approaches. Given the simplicity of the setup and the use of a self-hosted open-weight model (Qwen3.5-27B), the gap to classical methods is smaller than one might expect, and stronger frontier models may close it further.
Scaling the LLM from 0.8B to 27B is essential for unconstrained code editing but provides no advantage for fixed-HP optimization. Figure˜2 shows that 0.8B is insufficient for unconstrained code editing (Karpathy Agent Code: 0.9910 vs 0.9814 with 27B), while fixed-HP methods saw no benefit from scaling (Karpathy Agent 14 HPs: 0.9904 vs 0.9908). All LLM methods used open-weight Qwen3.5, and the gap between unconstrained code editing and classical methods is already small; stronger frontier models may further improve LLM-based methods, particularly for unconstrained code editing where model capability directly determines the quality of code modifications. A preliminary comparison with Gemini 2.5 Flash (Comanici et al., 2025) (Section˜A.2) shows no improvement over Qwen3.5-27B, though stronger or larger frontier models may yield different results.
5.3 Hybrid optimization: best of both worlds
Centaur outperformed all methods including CMA-ES alone by using the LLM on only 30% of trials. As described in Section˜4, the LLM receives CMA-ES’s full internal state ($\bm{\mu}$, $\sigma$, $\mathbf{C}$), the top-5 configurations, and the last 20 trials, and almost always overrides CMA-ES’s proposal (100% for 27B, 95% for 0.8B). Despite constituting only 30% of trials, LLM trials contributed 25% of incumbent improvements (Section˜A.4), confirming that the LLM provides complementary value beyond what CMA-ES finds alone. Beyond improving the mean, Centaur substantially reduces CMA-ES’s cross-seed variance: std drops from 0.0036 for CMA-ES alone to 0.0005 for Centaur, suggesting that the LLM stabilizes the optimizer by injecting domain knowledge that prevents unfavorable seeds from drifting. Notably, Centaur [0.8B] outperformed Centaur [27B] (0.9766 vs 0.9763), suggesting that a cheap LLM suffices when paired with a strong classical optimizer. This stands in contrast to unconstrained code editing, where the 0.8B model is insufficient: when the classical optimizer handles the search trajectory, the LLM’s role reduces to refining promising candidates, a task that does not require the same level of model capability as generating code edits from scratch. We ablate the LLM ratio $r$ in Section˜A.6. For the 0.8B model, r=0.2r{=}0.2 achieves the best single-seed result (0.9735), slightly outperforming the default r=0.3r{=}0.3 (0.9776), while r=0.8r{=}0.8 degrades to 0.9789. For the 27B model, r=0.5r{=}0.5 performs best (0.9745), but r=0.8r{=}0.8 collapses to 0.9903, worse than CMA-ES alone. This confirms that CMA-ES should retain majority control of the optimization trajectory: the LLM is most effective as an occasional informed perturbation, not as the primary search driver.
6 Conclusion
We benchmarked classical, LLM-based, and hybrid HPO methods on small-scale language model training under the same budgets and constraints. Within the fixed, constrained search space, classical HPO methods such as CMA-ES and TPE consistently outperformed LLM-based agents, with OOM avoidance emerging as a stronger predictor of performance than search diversity. Restricting LLMs to a fixed hyperparameter search space does not leverage their strengths; the one setting where a pure LLM method competes with classical approaches is unconstrained code editing (Karpathy, 2025a), where the agent directly modifies training source code and narrows the gap to classical methods substantially despite using only a self-hosted open-weight model. However, this requires sufficient model scale: the 0.8B model is insufficient for unconstrained code editing, while it suffices for hybrid optimization. Centaur achieved the best result in our experiments by using the LLM on only 30% of trials, preserving fast convergence from CMA-ES while benefiting from LLM-informed suggestions, with the 0.8B variant outperforming the 27B variant.
Our study evaluated a single task with open-weight models (Qwen3.5 0.8B and 27B). Given that unconstrained code editing already nearly matches classical methods with these models, stronger frontier models may shift the balance further in favor of LLM-based approaches. While we automated hyperparameter selection, the search space ranges still require manual specification, introducing some human priors. Future work could benchmark unconstrained code editing with frontier models to determine whether increased model capability allows LLM agents to match or surpass hybrid and classical methods. Additionally, exploring other classical optimizers as the hybrid base and pairing CMA-ES with a code-editing LLM agent could allow the search space to co-evolve with the optimization trajectory.
Appendix A Additional Results
A.1 Convergence by Trial Number
The main text reports convergence against cumulative training time, which is our primary comparison. Figures˜3 and 4 show convergence by trial number instead, measuring sample efficiency. These views can differ substantially because LLM-based methods spend additional time on inference between trials, compressing their wall-time curves even when they are competitive per trial.
Figure 3: Convergence by trial number (mean $\pm$ std across 3 seeds). Same methods as Figure˜1. Trial-number view shows sample efficiency rather than wall-clock cost.
Figure 4: 0.8B vs 27B by trial number (mean $\pm$ std across 3 seeds). Same methods as Figure˜2. Solid lines: 27B, dashed: 0.8B.
A.2 Frontier Model Comparison: Gemini 2.5 Flash
To test whether a stronger LLM optimizer changes the balance between classical and LLM-based methods, we ran Centaur, Karpathy Agent (Code), and LLAMBO (Paper) with Gemini 2.5 Flash [Comanici et al., 2025] as the LLM optimizer instead of Qwen3.5-27B. Figure˜5 compares the two over cumulative training time. Classical methods (TPE, Random) are shown as references. Within this window, Gemini 2.5 Flash does not outperform Qwen3.5-27B on this task, suggesting that simply swapping to a frontier model does not automatically close the gap to classical methods.
Figure 5: Gemini 2.5 Flash vs Qwen3.5-27B as LLM optimizer (cumulative training time). Solid: Gemini, dashed: Qwen3.5-27B. Same color per method.
Figure 6: Incumbent traces for classical and hybrid methods (wall-time, best seed). Grey dots: all trials. Colored dots: new incumbents. Staircase: best-so-far.
A.3 Incumbent Traces
Beyond aggregate convergence, we visualize per-method incumbent traces to reveal when and how each optimizer finds improvements. Figures˜6 and 7 show incumbent traces against cumulative training time (hours) for all methods. Grey dots are all trials, colored dots are new incumbents, and the staircase line is the best-so-far trajectory. Each panel shows the best seed for that method.
Figure 7: Incumbent traces for LLM-based methods (wall-time, best seed).
A.4 Qualitative Agent Behavior
The convergence plots show that Centaur improves over CMA-ES alone; we now examine how the LLM uses the shared optimizer state in practice. In Centaur seed 0, trial 136 produced a new incumbent (val_bpb=0.9837{}=0.9837):
  • CMA-ES suggested: WINDOW_PATTERN=LLLL, DEVICE_BATCH_SIZE=61, TOTAL_BATCH_SIZE=133143, SCALAR_LR=0.208
  • LLM overrode to: WINDOW_PATTERN=SSSS, DEVICE_BATCH_SIZE=64, TOTAL_BATCH_SIZE=131072, SCALAR_LR=0.3
Three overrides are notable: (1) Attention pattern ($LL\toSSSS$): CMA-ES has no domain knowledge about attention patterns. The LLM chose all-short attention, which is memory-efficient at the given depth (DEPTH=10{}=10). This is transformer-specific knowledge that CMA-ES cannot learn from scalar loss values. (2) Hardware-friendly rounding: the LLM chose power-of-2 batch sizes (64, 131072) instead of CMA-ES’s arbitrary values (61, 133143), aligning with GPU memory and tensor core constraints. (3) Learning rate: the LLM boosted SCALAR_LR from 0.208 to 0.3, closer to the regime where good configs cluster.
Overall in seed 0, 6 of 24 incumbent improvements came from LLM trials (25%), while LLM trials constituted 88 out of 275 total trials (32%, close to the 30% target ratio).
A.5 Diversity Metrics
We now define the diversity metrics reported in Table˜3. We compute all metrics on the 13 continuous HPs, excluding the categorical WINDOW_PATTERN. We normalize each HP to $[0,1]$ within its bounds and use only successful (non-OOM) trials.
  • Spread: mean standard deviation per HP across all trials. Higher values indicate more diverse sampling across each dimension.
  • Pairwise: mean L2 distance between all pairs of configurations. Higher values indicate configs are more different from each other.
  • Step: mean L2 distance between consecutive trials. Higher values indicate larger jumps between suggestions.
  • Cells: we discretized each HP into 5 equal-width bins and counted the number of unique 13-dimensional bin vectors across all trials. The theoretical maximum is $5^{13}\approx 1.2\times 10^{9}$; values in Table˜3 range from 29 to 805, indicating that all methods cover a small fraction of the search space.
A.6 Centaur LLM Ratio Ablation
We ablate the fraction of trials delegated to the LLM in Centaur. Table˜4 reports best val_bpb for each ratio on seed 0. The default ratio of r=0.3r{=}0.3 balances classical and LLM contributions; higher ratios degrade performance, particularly for the 27B model where r=0.8r{=}0.8 performs worse than CMA-ES alone. This confirms that CMA-ES needs to retain majority control of the optimization trajectory, with the LLM contributing occasional informed suggestions rather than dominating the search.
Table 4: Centaur LLM ratio ablation (seed 0). Best val_bpb for each ratio. Bold = best per column. The default ratio is r=0.3r{=}0.3.
LLM Ratio $r$ Centaur [0.8B] Centaur [27B]
0.1 0.9746 0.9752
0.2 0.9735 0.9753
0.3 (default) 0.9776 0.9756
0.5 0.9787 0.9745
0.8 0.9789 0.9903
Figure 8: Effect of LLM ratio on Centaur performance (seed 0). Higher ratios give the LLM more trials. CMA-ES alone (dotted) shown as reference. Too much LLM control degrades performance, especially for the 27B model at r=0.8r{=}0.8.
A.7 0.8B LLM Variant Results
The main text focuses on 27B variants; we now report the corresponding 0.8B results. Table˜5 reports diversity metrics for the 0.8B LLM variants of all LLM-based methods.
Table 5: Diversity analysis for 0.8B LLM variants (3 seeds each).
Method Trials Best val_bpb OOM% Spread Pairwise Step Cells
[…]
A.8 LLAMBO (Optuna) vs LLAMBO (Paper): Detailed Comparison
We include two LLAMBO variants because the OptunaHub port differs from the original paper in ways that substantially affect performance. Table˜6 summarizes the key implementation differences.
Table 6: Key differences between LLAMBO (Optuna) and LLAMBO (Paper).
Aspect Original Paper OptunaHub Port
Surrogate labels Continuous metric values Binary 0/1 (top 20% threshold)
Categorical HPs All HPs in LLM prompts Delegated to random sampling
Failed trials Visible to surrogate Hidden (TrialState.FAIL)